# Phase transitions and scale invariance in topological Anderson insulators

### v1

@@ Abstract
[…]
Figure 4: (a-b) Phase diagrams calculated using the local Chern marker $\mathcal{C}_{0}$ as a function of the BHZ mass $M$ and disorder strength $W$ for (a) $L=50$ nm and (b) 100 nm. The dashed white line marks the phase boundary between the trivial and topological phases (TI and TAI) obtained from the gapless condition $A/|\widetilde{M}|\ll L$. The dot-dashed line marks the boundary where $W$ matches the band width of the clean system, above which the Anderson localization phase takes place. (c) All data from panels (a) and (b), and additional data for a system of $L=150$ nm, fall into a universal curve for the trivial to TAI phase transition upon the scaling $(M,W,L)\rightarrow z=[W^{\mu}-W_{c}^{\mu}(M)]L/W^{\mu}_{\rm max}L_{\rm max}$.
The data shown in Figs. 4(a) and (b) for all $M$ and $W$, and additional data for $L=150$ nm (not shown), can be combined into a single universal function that characterizes the trivial to TAI phase transition shown in Fig. 4(c). The scaling used in this case is a direct extension of the one presented in Eqs. (5) and (6), where now $W_{c}\rightarrow W_{c}(M)$ depends upon $M$. To obtain an expression for $W_{c}(M)$ and to understand the value $\mu\approx 2$ of the critical parameters, let us first consider the clean regime ($W=0$), to find the scaling law for the trivial to TI phase transition as a function of $M$. Here, we consider an argument similar to the one used in Ref. [20] for the Haldane model, as follows. In the topological regime, the only relevant length scales are $\xi=A/|M|$ and $L$, as introduced above. Therefore, one can expect that the topological phase transition scaling with the system size $L$ should be a universal function of $L/\xi\propto ML$ (assuming $A$ is constant). Thus, we justify that
$\mathcal{C}_{0}(M,L)\equiv\mathcal{C}_{0}(ML),$ (8)
@@
Conclusions.– We have investigated the scale invariance of the local Chern marker $\mathcal{C}_{0}$ applied to topological Anderson insulators (TAI). The smooth profile of $\mathcal{C}_{0}$ throughout the phase transition allows us to characterize the scale invariance, obtain critical parameters, and universal scaling functions. Our analysis uses the disordered BHZ model as a canonical example of TI and TAI phases. The simplicity of the model allows for an analytical justification of the obtained numerical results. We like to stress that the disordered Haldane [42] and Kane-Mele [43] models exhibit an identical scaling and universal behavior, as we have shown in the Supplemental Material [44]. This observation supports the statement that our findings are quite generic. Indeed, as long as the first-order Born approximation remains sufficiently accurate, we expect that the universal function scales with $\sim W^{2}L$, while deviations from this behavior should occur only beyond the validity range of the 1BA. We believe that our findings shed light on an approach to investigating scaling properties of topological phase transitions for disordered and amorphous systems.
Note added: shortly before we finished this Letter, Ref. [52] - that has a large overlap with our study - appeared in the arXiv.
Acknowledgments.– This work was supported by the Brazilian funding agencies CNPq, CAPES, FAPERJ, and FAPEMIG (grant PPM-00798-18). The authors acknowledge useful conversation with C. Beenakker, A. Akhmerov, and C. P. Orth.

### v2

@@ Abstract
Here and in what follows we use that the first Born approximation gives $\alpha(M)\approx\alpha_{0}$ [59]. In Figs. 4(a) and 4(b) the dashed lines mark the phase boundary given by $W_{\rm gapless}(M)$ for an optimal $\alpha_{0}\approx 1/3700$ meV-1, which matches remarkably well the phase boundary between the trivial and topological phases (TI and TAI).
Figure 4: (a, b) Phase diagrams calculated using the local Chern marker $\mathcal{C}_{0}$ as a function of the BHZ mass $M$ and the disorder strength $W$ for (a) $L=50$ nm and (b) 100 nm. In all cases, we consider 200 disorder realizations and 1000 Chebyshev moments on the KPM expansion. The dashed white line marks the phase boundary between the trivial and topological phases (TI and TAI) obtained from the gapless condition $A/|\widetilde{M}|\ll L$. The dot-dashed line marks the boundary where $W$ matches the bandwidth of the clean system, above which the Anderson localization phase takes place. (c) All data from panels (a) and (b), and additional data for a system of $L=150$ nm, fall into a universal curve for the trivial to TAI phase transition upon the scaling $(M,W,L)\rightarrow z=[W^{\mu}-W_{c}^{\mu}(M)]L/W^{\mu}_{\rm max}L_{\rm max}$.
The data shown in Figs. 4(a) and 4(b) for all $M$ and $W$, and additional data for $L=150$ nm (not shown), can be combined into a single universal function that characterizes the trivial to TAI phase transition shown in Fig. 4(c). The scaling used in this case is a direct extension of the one presented in Eqs. (5) and (6), where now $W_{c}\rightarrow W_{c}(M)$ depends upon $M$. To obtain an expression for $W_{c}(M)$ and to understand the value $\mu\approx 2$ of the critical parameters, let us first consider the clean regime ($W=0$), to find the scaling law for the trivial to TI phase transition as a function of $M$. Here, we consider an argument similar to the one used in Ref. [21] for the Haldane model, as follows. In the topological regime, the only relevant length scales are $\xi=A/|M|$ and $L$, as introduced above. Therefore, one can expect that the topological phase transition scaling with the system size $L$ should be a universal function of $L/\xi\propto ML$ (assuming $A$ is constant). Thus, we justify that
$\mathcal{C}_{0}(M,L)\equiv\mathcal{C}_{0}(ML).$ (8)
@@
Conclusions.– We have investigated the scale invariance of the local Chern marker $\mathcal{C}_{0}$ applied to TAIs. The smooth profile of $\mathcal{C}_{0}$ throughout the phase transition allows us to characterize the scale invariance and obtain critical parameters, and universal scaling functions. Our analysis uses the disordered BHZ model as a canonical example of TI and TAI phases. The simplicity of the model allows for an analytical justification of the obtained numerical results. We like to stress that the disordered Haldane [57] and Kane-Mele [58] models exhibit an identical scaling and universal behavior, as we have shown in the Supplemental Material [59]. This observation supports the statement that our findings are quite generic. Indeed, as long as the first-order Born approximation remains sufficiently accurate, we expect that the universal function will scale with $\sim W^{2}L$, while deviations from this behavior should occur only beyond the validity range of the 1BA. We believe that our findings shed light on an approach to investigating scaling properties of topological phase transitions for disordered and amorphous systems.
Note added in proof. The independent study [72] has considerable overlap with our work.
Acknowledgments.– This work was supported by the Brazilian funding agencies CNPq, CAPES, FAPERJ, and FAPEMIG (Grant No. PPM-00798-18). The authors acknowledge useful conversations with C. Beenakker, A. Akhmerov, and C. P. Orth.